OpenStudy (anonymous):

square root of 1/121

OpenStudy (anonymous):

\[\LARGE \sqrt{\frac1{121}}\] \[\Large \sqrt{\frac{m}n}= \frac{\sqrt{m}}{\sqrt n}\]

OpenStudy (anonymous):

simplfy it

OpenStudy (anonymous):

That's your job, isn't it? :3 I've already given you a pretty strong hint...

OpenStudy (anonymous):

kk